Use logarithms to solve the given equation. (Rou 42x + 1 = 27 x =
Answer #1

42x + 1 = 27

Taking logarithm on both sides we get,

log (42X + 1) = log (27)

Or, (2X + 1) log (4) = log (27)

Or, 2X + 1 = (log 27)/(log 4)

Or, 2X + 1 = 2.377443751

Or, 2X = 1.377443751

Or, X = 0.6887 = 0.69 (approximate)

Answer: 0.69